OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (−6, 3). y = 3/2x − 1 y = 3/2x + 3 y = −2/3x − 1 y = −2/3x + 3

OpenStudy (anonymous):

what do you know about slope of perpendicular line?

OpenStudy (anonymous):

perpendicular slope will be -1

OpenStudy (anonymous):

using the perpendicular slope and point (-6,3) the line will be. . .

OpenStudy (anonymous):

the answer is not found on the choices. .

OpenStudy (anonymous):

using point-slope form, given the perpenducular lope and the given point, the equation of the line will be y= -x -3

OpenStudy (mysesshou):

point (−6, 3). y = x + 4 m=-1 y+6 = -1*(x-3) y= -x+3-6 y=-x-3 agree with jerwyn gayo also easily seen as original slope is 1, so perpendicular slope is -1, and none of the answers have -1 as the slope.